# Analysis of Hessian Scaling for Local and Global Costs in Variational Quantum Algorithms

### v1

@@ 4.2.2 VQE with Transverse Field Ising Model
To validate our theoretical findings in a realistic variational quantum eigensolver (VQE) setting [kandala2017hardware], we replace the toy objective with the Hamiltonian of the 1D Transverse Field Ising Model (TFIM) [pfeuty1970one] under periodic boundary conditions:
@@
Figure 3: Hessian-entry variance vs. system size in a VQE task (TFIM). Comparison of the variances of diagonal (left) and off-diagonal (right) Hessian entries between the VQE local cost (blue squares) and the global parity cost (red circles). The dashed lines represent linear regression fits.
@@ 4.2.3 Dependence on depth
We next examine how Hessian-entry fluctuations depend on circuit depth $L$, which provides a numerical view of how increasing operator spreading affects second-order resolvability. Figure 4 reports empirical variances of representative diagonal and off-diagonal entries at fixed system size $n=16$ over the depth range $L\in\{1,\ldots,12\}$.
For $C_{\mathrm{global}}$ (red markers), both $\operatorname{Var}_{\rho}(H_{jj})$ and $\operatorname{Var}_{\rho}(H_{jk})$ are already on the order of $10^{-5}$ by depth $L=2$ and exhibit only weak additional dependence on $L$ over the explored range. This behavior is consistent with the global-support regime: backward lightcones overlap extensively even at shallow depth, so cancellations can suppress entrywise curvature fluctuations early, leaving little variation to resolve at the scale of our reported statistics.
@@ 4.2.4 Robustness against measurement shot noise
In Fig. 5(a), the measured curves are well fit by this model across the tested shot range. The separation of a constant term and a $1/N$ term matches the variance decomposition in Corollary 3.9. Comparing objectives, the fitted parameters reflect the distinct $n$-scalings predicted by Theorem 3.3: the global objective exhibits rapidly shrinking intrinsic fluctuations with $n$ but correspondingly stringent resolution demands, whereas the termwise local averaged objective follows a polynomial regime.
To summarize the practical implications, Fig. 5(b) reports an empirical shot requirement $N(\varepsilon)$ for a prescribed tolerance $\varepsilon$, defined as the smallest $N$ such that the fitted model satisfies $\sqrt{\operatorname{Var}_{\rho,\mathrm{sh}}(\widehat{H}_{jk})}\leq\varepsilon$. The resulting curves show that achieving a fixed absolute tolerance becomes rapidly more expensive with $n$ for the global objective, while the termwise local averaged objective remains substantially less demanding, consistent with the different $n$-scalings of $\operatorname{Var}_{\rho}(H_{jk})$ in Theorem 3.3.
Figure 5: Impact of finite measurement on Hessian estimation. Same Hessian entry and initialization protocol under a global objective and a termwise local averaged objective. (Left) Total estimator variance $\operatorname{Var}_{\rho,\mathrm{sh}}(\widehat{H}_{jk})$ versus shots NN. (Right) Empirical shot complexity $N(\varepsilon)$ required to reach $\sqrt{\operatorname{Var}_{\rho,\mathrm{sh}}(\widehat{H}_{jk})}\leq\varepsilon$ versus $n$.
@@ 4.3 Spectral structure
While our theory controls entrywise curvature fluctuations, curvature-aware procedures also depend on how curvature is distributed across parameter-space directions. We therefore report complementary spectral diagnostics based on the empirical eigenvalue distribution of the Hessian at random initialization. These results are presented as numerical diagnostics and are not interpreted as conditioning guarantees.
@@
Fig. 6 shows histogram estimates (log-density scale) of Hessian eigenvalues for global and termwise local averaged objectives at depth $L=4$ and increasing system size. As $n$ grows, the spectrum associated with the global objective becomes increasingly concentrated near the origin, whereas the spectrum for the local objective remains comparatively broader over the same range. This qualitative contrast mirrors the entrywise variance behavior: global objectives exhibit a rapidly shrinking initialization-scale curvature distribution, while locality delays such concentration for termwise averaged costs.
@@
Fig. 7 (Left) shows that $\lambda_{\mathrm{RMS}}$ decreases with $n$ for both objectives, consistent with an overall contraction of typical curvature magnitudes at initialization. Fig. 7 (Right) shows that, at the same numerical tolerance, the global objective exhibits a substantially larger near-zero fraction than the local averaged objective. Interpreted at fixed resolution, this indicates more directions that appear numerically flat at initialization for the global objective than for the termwise averaged local objective.
@@
Taken together with the entrywise variance results in § 4, these spectral diagnostics provide geometric context for the observed variance scalings.
4.4 Budget-limited manifestations
The variance scalings in § 3 have direct implications for what curvature information is resolvable under finite evaluation complexity at initialization. To illustrate this point, we report simple budget-matched optimization trajectories for global and termwise local averaged objectives.
We consider $(n,L)=(18,8)$ and compare stochastic gradient descent (SGD) [amari1993backpropagation] with a representative geometry-aware first-order method, the quantum natural gradient (QNG) [stokes2020quantum], using matched iteration complexity and fixed step-size schedules. QNG is included only as a preconditioned first-order baseline; the purpose is not an optimizer comparison but an illustration of budget-limited signal resolution near initialization.
Fig. 8 shows mean cost trajectories with one-standard-deviation bands over random initializations. For the global objective $C_{\mathrm{global}}$ (red), both SGD and QNG exhibit negligible progress at the displayed resolution, and trajectories concentrate tightly across seeds. In contrast, the termwise averaged local objective $C_{\mathrm{local}}$ (blue) decreases reliably and shows appreciable seed-to-seed variability under the same budget. These observations are consistent with the entrywise variance regimes in § 3 and with the finite-budget interpretation of Corollary 3.9: under a fixed estimator resolution, global objectives quickly lose resolvable descent directions at initialization, whereas locality preserves observable signals over the same depth and budget.
Figure 8: Budget-limited optimization trajectories at $(n,L)=(18,8)$. Mean cost versus iteration for $C_{\mathrm{global}}$ (red) and $C_{\mathrm{local}}$ (blue): SGD (left) and QNG (right). Shading: $\pm 1$ standard deviation over random initializations.
@@ 5 Conclusion
In this study, we have systematically analyzed the second-order trainability of variational quantum algorithms by quantifying the initialization scaling of Hessian-entry fluctuations. By leveraging exact parameter-shift rules to map entrywise variances to finite covariance quadratic forms, we established a fundamental dichotomy in the statistical resolvability of curvature information. For global objectives, the Hessian inherits exponential concentration, implying that resolving curvature requires a prohibitive exponential sampling budget. Conversely, for termwise $k$-local objectives, the variance decay is polynomial and governed by the causal lightcone volume. These findings delineate the precise boundary where second-order information remains computationally accessible, a distinction that is rigorously supported by our numerical experiments across varying system configurations.
Several avenues remain for future research. Theoretically, identifying conditions under which entrywise resolvability implies favorable matrix conditioning would strengthen the link to optimizer convergence. Practically, extending this variance-based framework to realistic noise models is essential for NISQ applications. Broadening the scope, our findings on derivative scaling have direct implications for advanced computational paradigms that rely on higher-order information, including quantum multiscale modeling [chen2022qm, deiml2025quantum, braun2025higher] and quantum machine learning [febrianto2025quantum, liu2025beyond]. Understanding the resolvability limits in these contexts will be crucial for developing robust quantum algorithms for complex scientific computing tasks.
@@ A.1 Covariance–quadratic representations from parameter shifts
$H_{jj}=\frac{1}{4}\big(C_{+}-2C_{0}+C_{-}\big)=\sum_{\alpha\in\{+,0,-\}}a_{\alpha}\,C_{\alpha},\qquad(a_{+},a_{0},a_{-})=\tfrac{1}{4}(1,-2,1).$
Using $\operatorname{Var}(\sum_{\alpha}a_{\alpha}C_{\alpha})=\sum_{\alpha,\beta}a_{\alpha}a_{\beta}\,\operatorname{Cov}(C_{\alpha},C_{\beta})$ yields
$\operatorname{Var}(H_{jj})=\sum_{\alpha,\beta\in\{+,0,-\}}a_{\alpha}a_{\beta}\,\operatorname{Cov}(C_{\alpha},C_{\beta})=a^{\top}\Sigma a,$
@@
$-4\operatorname{Cov}(C_{+},C_{0})-4\operatorname{Cov}(C_{-},C_{0})\Big).$
which is an explicit linear combination of variances and covariances among $\{C_{+},C_{0},C_{-}\}$, as stated in Lemma 2.3.
@@ A.1.2 Off-diagonal entries
Let $c:=(C_{++},C_{+-},C_{-+},C_{--})^{\top}$ and $s:=(1,-1,-1,1)^{\top}$, and set $\Sigma:=\operatorname{Cov}(c)\in\mathbb{R}^{4\times 4}$. Then we have $\operatorname{Var}(H_{jk})=\frac{1}{16}s^{\top}\Sigma s$. Equivalently,
(23) $\operatorname{Var}(H_{jk})=\frac{1}{16}\sum_{\sigma,\tau\in\{+,-\}}\sum_{\sigma^{\prime},\tau^{\prime}\in\{+,-\}}s_{\sigma\tau}\,s_{\sigma^{\prime}\tau^{\prime}}\operatorname{Cov}(C_{\sigma\tau},C_{\sigma^{\prime}\tau^{\prime}}),$
where $s_{++}=s_{--}=1$ and $s_{+-}=s_{-+}=-1$.
@@
$X(\bm{\theta})=\sum_{\ell=1}^{m}a_{\ell}C(\bm{\theta}\oplus s_{\ell}),\qquad m<\infty,$
@@
$\operatorname{Var}_{\rho}(X)=\sum_{\ell,\ell^{\prime}=1}^{m}a_{\ell}a_{\ell^{\prime}}\operatorname{Cov}_{\rho}\big(C(\bm{\theta}\oplus s_{\ell}),C(\bm{\theta}\oplus s_{\ell^{\prime}})\big).$
In our setting, the relevant shift sets have constant size $m=O(1)$ determined only by the chosen parameter-shift rule.
@@ A.3 A dependency-graph variance bound
Step 0: Finite-shift covariance representations
By the parameter-shift identities (Assumption 2.1), each Hessian entry $H_{jk}$ can be written as a finite linear combination of shifted costs:
$H_{jk}=\sum_{s\in\mathcal{S}_{ab}}a_{s}\,C(\bm{\theta}\oplus s),\qquad|\mathcal{S}_{ab}|=O(1).$
Hence
(29) $\operatorname{Var}(H_{jk})=\sum_{s,s^{\prime}\in\mathcal{S}_{ab}}a_{s}a_{s^{\prime}}\,\operatorname{Cov}\!\big(C(\bm{\theta}\oplus s),\,C(\bm{\theta}\oplus s^{\prime})\big).$
For diagonal entries one may take $\mathcal{S}_{jj}=\{+\pi e_{j},0,-\pi e_{j}\}$ and recover (13); for off-diagonal entries one may take $\mathcal{S}_{jk}=\{\pm\frac{\pi}{2}e_{j}\pm\frac{\pi}{2}e_{k}\}$ and recover Lemma A.2. It therefore suffices to control covariances between shifted costs.
Throughout, Assumption 2.4 implies $|C(\cdot)|\leq 1$, and thus $|\operatorname{Cov}(\cdot,\cdot)|\leq 1$.
@@
Substituting this uniform bound into (29) and using that $|\mathcal{S}_{ab}|=O(1)$ gives
@@
Step 2: Termwise local averaged objectives (polynomial scaling)
Assume the objective is of the termwise local averaged form
@@
Finally, substitute (31) into the covariance–quadratic representation (29). Since $|\mathcal{S}_{ab}|=O(1)$ and the coefficients $\{a_{s}\}$ are constants determined by the shift rule, we obtain
@@ B.3 From entry-wise second moments to matrix norm bounds
We state in Lemma 3.7 the corresponding matrix norm implications in the main text; we provide a complete proof here for completeness.
Lemma B.5 (From entry-wise second moments to Frobenius and spectral norm bounds).
(37) $\mathbb{E}\|H\|_{F}^{2}=\sum_{j,k=1}^{M}\mathbb{E}[H_{jk}^{2}]=\sum_{j,k=1}^{M}\Big(\operatorname{Var}(H_{jk})+(\mathbb{E}[H_{jk}])^{2}\Big).$
Moreover, since $\|H\|_{2}\leq\|H\|_{F}$ pointwise, we have the second-moment bound
(38) $\mathbb{E}\|H\|_{2}^{2}\leq\mathbb{E}\|H\|_{F}^{2},$
and by Jensen’s inequality,
(39) $\mathbb{E}\|H\|_{2}\leq\mathbb{E}\|H\|_{F}\leq\sqrt{\mathbb{E}\|H\|_{F}^{2}}.$
In particular, if $\operatorname{Var}(H_{jk})\leq v_{n}$ for all $j,k$ and $|\mathbb{E}H_{jk}|\leq\mu_{n}$ for all $j,k$, then
(40) $\mathbb{E}\|H\|_{F}^{2}\leq M^{2}(v_{n}+\mu_{n}^{2}),\qquad\mathbb{E}\|H\|_{2}^{2}\leq M^{2}(v_{n}+\mu_{n}^{2}),\qquad\mathbb{E}\|H\|_{2}\leq M\sqrt{v_{n}+\mu_{n}^{2}}.$
Corollary B.6 (Curvature magnitude scale (entry-wise to global norms)).
Under the hypotheses of Theorem 3.3, suppose additionally that there exists a constant $c>0$ independent of $n$ such that for all $j,k$,
$|\mathbb{E}[H_{jk}]|\leq\mu_{n}\qquad\text{and}\qquad\mu_{n}^{2}\leq c\,v_{n},$
where $v_{n}$ denotes the entry-wise variance bound from Theorem 3.3. Then
$\mathbb{E}\|H\|_{F}^{2}\lesssim M^{2}v_{n},\qquad\mathbb{E}\|H\|_{2}^{2}\lesssim M^{2}v_{n},\qquad\mathbb{E}\|H\|_{2}\lesssim M\sqrt{v_{n}}.$
These bounds quantify a global magnitude scale for curvature at initialization (and hence for the typical size of Hessian–vector products), but they do not, by themselves, determine conditioning, $\lambda_{\min}(H)$, or near-singularity without further spectral information.

### v2

@@ 4.2.2 VQE with Transverse Field Ising Model
To validate our theoretical findings in a realistic variational quantum eigensolver (VQE) setting [kandala2017hardware], we replace the toy objective (cf. § 4.1) with the Hamiltonian of the 1D Transverse Field Ising Model (TFIM) [pfeuty1970one] under periodic boundary conditions:
@@
Figure 3: Hessian-entry variance vs. system size in a VQE task (TFIM). Comparison of the variances of diagonal (left) and off-diagonal (right) Hessian entries between the VQE local cost (blue squares) and the global parity cost (red circles).
@@ 4.2.3 Dependence on depth
We next examine how Hessian-entry fluctuations depend on circuit depth $L$, which provides a numerical view of how increasing operator spreading affects second-order resolvability. Fig. 4 reports empirical variances of representative diagonal and off-diagonal entries at fixed system size $n=16$ over the depth range $L\in\{1,\ldots,12\}$.
For $C_{\mathrm{global}}$ (red markers), both $\operatorname{Var}_{\rho}(H_{jj})$ and $\operatorname{Var}_{\rho}(H_{jk})$ are already on the order of $10^{-5}$ by depth $L=2$ and exhibit only weak additional dependence on $L$ over the explored range. This behavior is consistent with the global-support regime: backward lightcones overlap extensively even at shallow depth, so cancellations can suppress entry-wise curvature fluctuations early, leaving little variation to resolve at the scale of our reported statistics.
@@ 4.2.4 Robustness against measurement shot noise
$\operatorname{Var}_{\rho,\mathrm{sh}}(\widehat{H}_{jk})=\operatorname{Var}_{\rho}(H_{jk})+\mathbb{E}_{\bm{\theta}\sim\rho}\left[\operatorname{Var}_{\mathrm{sh}}(\widehat{H}_{jk}\mid\bm{\theta})\right].$
Accordingly, we fit the two-term model $\operatorname{Var}_{\rho,\mathrm{sh}}(\widehat{H}_{jk})\approx a+b/N$, where $a$ estimates the intrinsic initialization variance up to shift-rule constants.
In Fig. 5(a), the data exhibit an $N$-independent floor and a $1/N$ decay, consistent with this model. Figure 5(b) reports an absolute-tolerance requirement $N(\varepsilon):=\min\{N:\sqrt{\operatorname{Var}_{\rho,\mathrm{sh}}(\widehat{H}_{jk})}\leq\varepsilon\}$ with $\varepsilon=0.05$; once $a\ll\varepsilon^{2}$, this criterion becomes shot-noise limited and $N(\varepsilon)\approx b/\varepsilon^{2}$, so the curves can saturate (or decrease) as $n$ increases. The scaling relevant to resolvability at initialization is instead captured by $a$ (equivalently, $\operatorname{Var}_{\rho}(H_{jk})$ up to shift-rule constants) and the resolution criterion in Definition 3.1 (Corollary 3.12), which ties the required shots to $\operatorname{Var}_{\rho}(H_{jk})^{-1}$.
Figure 5: Finite-shot effects on Hessian-entry estimation. (Left) $\operatorname{Var}_{\rho,\mathrm{sh}}(\widehat{H}_{jk})$ versus shots NN. (Right) Absolute-tolerance requirement $N(\varepsilon)$ with $\varepsilon=0.05$ versus $n$.
@@ 4.3 Spectral structure
While our theory controls entry-wise curvature fluctuations, curvature-aware procedures also depend on how curvature is distributed across parameter-space directions. We therefore report complementary spectral diagnostics based on the empirical eigenvalue distribution of the Hessian at random initialization. These results are presented as numerical diagnostics and are not interpreted as conditioning guarantees.
@@
Fig. 6 shows histogram estimates (log-density scale) of Hessian eigenvalues for global and term-wise local averaged objectives at depth $L=4$ and increasing system size. As $n$ grows, the spectrum associated with the global objective becomes increasingly concentrated near the origin, whereas the spectrum for the local objective remains comparatively broader over the same range. This qualitative contrast mirrors the entry-wise variance behavior: global objectives exhibit a rapidly shrinking initialization-scale curvature distribution, while locality delays such concentration for term-wise averaged costs.
@@
Fig. 7 (Left) shows that $\lambda_{\mathrm{RMS}}$ decreases with $n$ for both objectives, consistent with an overall contraction of typical curvature magnitudes at initialization. Fig. 7 (Right) shows that, at the same numerical tolerance, the global objective exhibits a substantially larger near-zero fraction than the local averaged objective. Interpreted at fixed resolution, this indicates more directions that appear numerically flat at initialization for the global objective than for the term-wise averaged local objective.
@@
4.4 Finite-shot behavior near initialization
The variance scalings in § 3 suggest that, under finite-shot sampling, global and term-wise local objectives can exhibit markedly different signal-to-noise behavior near random initialization. To illustrate this effect, we report matched optimization trajectories for a global objective and a term-wise local averaged objective under the same measurement protocol.
We fix $(n,L)=(18,8)$ and run stochastic gradient descent (SGD) [amari1993backpropagation] and the quantum natural gradient (QNG) [stokes2020quantum]. QNG is included only as a geometry-aware first-order baseline; the goal is not an optimizer comparison, but to probe how finite sampling affects observable progress near initialization.
Fig. 8 reports the mean cost trajectories with $\pm 1$ standard deviation bands over random initializations. For the global objective $C_{\mathrm{global}}$ (red), both methods exhibit trajectories that are nearly indistinguishable across seeds and show little change over the plotted horizon at the chosen sampling level. However, the term-wise local averaged objective $C_{\mathrm{local}}$ (blue) decreases consistently and displays visible seed-to-seed variability under the same protocol. These trends are consistent with § 3: for global objectives, $\operatorname{Var}_{\rho}(H_{jk})$ decays exponentially in $n$, so at a fixed shot level the corresponding entry-wise fluctuations become difficult to resolve; for term-wise local objectives, the variance bounds are polynomial, so the same sampling protocol can still reveal measurable variation across initializations and iterations.
Figure 8: Finite-shot optimization trajectories at $(n,L)=(18,8)$. Mean cost versus iteration for $C_{\mathrm{global}}$ (red) and $C_{\mathrm{local}}$ (blue): SGD (left) and QNG (right). Shading: $\pm 1$ standard deviation over random initializations.
@@ 5 Conclusion
In this study, we quantified the scaling of entry-wise Hessian variances in variational quantum algorithms. Using exact second-order parameter-shift identities, we expressed each Hessian entry as a constant-size linear combination of shifted objective evaluations, thereby yielding a finite covariance–quadratic representation for $\operatorname{Var}_{\rho}(H_{jk})$. This framework leads to two distinct scaling behaviors for the resolvability of Hessian entries under finite-shot sampling. For global objectives satisfying a standard second-moment concentration condition, $\operatorname{Var}_{\rho}(H_{jk})$ decays exponentially in the number of qubits, and the corresponding resolution cost for Hessian-entry estimation grows exponentially with $n$. For term-wise $k$-local objectives in bounded-depth circuits, we obtain polynomial variance bounds governed by backward-lightcone growth on the interaction graph, which implies polynomial resolution cost whenever the lightcone remains sub-extensive. These conclusions are supported by numerical experiments across system sizes, circuit depths, and graph families.
Several avenues remain for future research. On the theory side, identifying conditions under which entry-wise resolvability correlates with informative spectral structure (e.g., conditioning or extremal eigenvalues) would clarify implications for Newton-type and preconditioned methods. On the practical side, extending the variance-based analysis to realistic noise models, including device noise and readout errors, is important for NISQ-relevant settings. More broadly, our results on derivative scaling inform higher-order quantum workflows in scientific computing, including quantum multiscale modeling [braun2025higher, chen2022qm, deiml2025quantum] and quantum machine learning [febrianto2025quantum, liu2025beyond]. Understanding resolvability limits in these contexts will be important for designing robust algorithms that rely on higher-order information.
@@ A.1 Covariance–quadratic representations from parameter shifts
$H_{jj}=\frac{1}{4}\big(C_{+}-2C_{0}+C_{-}\big)=\sum_{\alpha\in\{+,0,-\}}w_{\alpha}\,C_{\alpha},\qquad(w_{+},w_{0},w_{-})=\tfrac{1}{4}(1,-2,1).$
Using $\operatorname{Var}(\sum_{\alpha}w_{\alpha}C_{\alpha})=\sum_{\alpha,\beta}w_{\alpha}w_{\beta}\,\operatorname{Cov}(C_{\alpha},C_{\beta})$ yields
$\operatorname{Var}(H_{jj})=\sum_{\alpha,\beta\in\{+,0,-\}}w_{\alpha}w_{\beta}\,\operatorname{Cov}(C_{\alpha},C_{\beta})=w^{\top}\Sigma w,$
@@
$-4\operatorname{Cov}(C_{+},C_{0})-4\operatorname{Cov}(C_{-},C_{0})\Big),$
an explicit linear combination of variances and covariances among $\{C_{+},C_{0},C_{-}\}$, as stated in Lemma 2.3.
@@ A.1.2 Off-diagonal entries
Let $c:=(C_{++},C_{+-},C_{-+},C_{--})^{\top}$ and $w:=(1,-1,-1,1)^{\top}$, and set $\Sigma:=\operatorname{Cov}(c)\in\mathbb{R}^{4\times 4}$. Then $\operatorname{Var}(H_{jk})=\frac{1}{16}w^{\top}\Sigma w$. Equivalently,
(21) $\operatorname{Var}(H_{jk})=\frac{1}{16}\sum_{\sigma,\tau\in\{+,-\}}\sum_{\sigma^{\prime},\tau^{\prime}\in\{+,-\}}w_{\sigma\tau}\,w_{\sigma^{\prime}\tau^{\prime}}\operatorname{Cov}(C_{\sigma\tau},C_{\sigma^{\prime}\tau^{\prime}}),$
where $w_{++}=w_{--}=1$ and $w_{+-}=w_{-+}=-1$.
@@
$X(\bm{\theta})=\sum_{\ell=1}^{m}w_{\ell}C(\bm{\theta}\oplus s_{\ell}),\qquad m<\infty,$
@@
$\operatorname{Var}_{\rho}(X)=\sum_{\ell,\ell^{\prime}=1}^{m}w_{\ell}w_{\ell^{\prime}}\operatorname{Cov}_{\rho}\big(C(\bm{\theta}\oplus s_{\ell}),C(\bm{\theta}\oplus s_{\ell^{\prime}})\big).$
In our setting, the relevant shift sets have constant size $m=O(1)$ determined only by the chosen parameter-shift rule. This is the general form of (12).
@@ A.3 A dependency-graph variance bound
Substituting this uniform bound into (2.3) and using that $|\mathcal{S}_{ab}|=O(1)$ gives
@@
Step 2: Term-wise local averaged objectives (polynomial scaling)
Assume the objective is of the term-wise local averaged form
@@
Finally, substitute (28) into the covariance–quadratic representation (2.3). Since $|\mathcal{S}_{ab}|=O(1)$ and the coefficients $\{a_{s}\}$ are constants determined by the shift rule, we obtain
@@ B.3 From entry-wise second moments to matrix norm bounds
We provide a proof of Lemma 3.9 stated in the main text.
Proof B.5 (Proof of Lemma 3.9).
Let $H\in\mathbb{R}^{M\times M}$ be a random matrix with finite second moments. By definition, $\|H\|_{F}^{2}=\sum_{j,k=1}^{M}H_{jk}^{2}$, and taking expectations gives
$\mathbb{E}\|H\|_{F}^{2}=\sum_{j,k=1}^{M}\mathbb{E}[H_{jk}^{2}]=\sum_{j,k=1}^{M}\Big(\operatorname{Var}(H_{jk})+(\mathbb{E}[H_{jk}])^{2}\Big).$
Since $\|H\|_{2}\leq\|H\|_{F}$ pointwise, we have $\mathbb{E}\|H\|_{2}^{2}\leq\mathbb{E}\|H\|_{F}^{2}$. Moreover, Jensen’s inequality yields
$\mathbb{E}\|H\|_{2}\leq\mathbb{E}\|H\|_{F}\leq\sqrt{\mathbb{E}\|H\|_{F}^{2}}.$
If $\operatorname{Var}(H_{jk})\leq v_{n}$ and $\big|\mathbb{E}[H_{jk}]\big|\leq\mu_{n}$ for all $j,k$, then
$\mathbb{E}\|H\|_{F}^{2}=\sum_{j,k=1}^{M}\Big(\operatorname{Var}(H_{jk})+\big(\mathbb{E}[H_{jk}]\big)^{2}\Big)\leq M^{2}(v_{n}+\mu_{n}^{2}),$
and the spectral norm bounds above give
$\mathbb{E}\|H\|_{2}^{2}\leq M^{2}(v_{n}+\mu_{n}^{2}),\qquad\mathbb{E}\|H\|_{2}\leq M\sqrt{v_{n}+\mu_{n}^{2}}.$
This completes the proof.